OpenStudy (anonymous):

limit x tends to 0 for f(x)= (cos(sinx)-cosx)/x^2

OpenStudy (anonymous):

I've tried several things, transformation formulae, series expansions etc etc. Very stuck.

OpenStudy (jhannybean):

\[\large \lim_{x\rightarrow 0 } \frac{\cos(\sin(x))-\cos(x)}{x^2}\]

OpenStudy (anonymous):

On mobile, so thanks for the formatting.

OpenStudy (astrophysics):

Split it into two limits, one should require squeeze theorem I believe

OpenStudy (anonymous):

Squeeze theorem doesn't make sense if you have a x^2 term I thought?

OpenStudy (jhannybean):

No, it does.

ganeshie8 (ganeshie8):

we cannot split the limit here because the individual limits don't exist

ganeshie8 (ganeshie8):

have u tried taylor ? looks the numerator is \(\mathcal{O}(x^3)\)

OpenStudy (astrophysics):

I think you maybe right no we can just use L'hopital's rule

OpenStudy (astrophysics):

0/0

imqwerty (imqwerty):

what i did - 1) simplified cosA+cosB 2)took the constants out ...... only constant which ws there ws -2 :P 3)applied L hospitality nd then i got (-2)x0 =0 :)

OpenStudy (astrophysics):

looks good @qwerty!

imqwerty (imqwerty):

:) thanks @Astrophysics

OpenStudy (jhannybean):

cos(A) \(\color{red}{\text{+}}\) cos(B)?

OpenStudy (empty):

Yeah I did L'H twice and got 0 as well

OpenStudy (astrophysics):

Yup did same thing haha

OpenStudy (anonymous):

Oh well, we're required to solve without L'Hospital. So anyone got any clues on that.

OpenStudy (empty):

Any other restrictions we should know about...?

OpenStudy (empty):

If you can use series expansions then saying you can't use L'H is practically meaningless imo

OpenStudy (anonymous):

And I don't know what or how Taylor Series are... No other restrictions.

ganeshie8 (ganeshie8):

series expansion works nicely

ganeshie8 (ganeshie8):

just show that the degree of numerator is at least 3 and you're done

OpenStudy (astrophysics):

Niceee

OpenStudy (anonymous):

Okay? @ganeshie8 and that's proof cause?

ganeshie8 (ganeshie8):

so are you allowed to use below ? \[\sin(x)=x-\frac{x^3}{3!} + \frac{x^5}{5!}\mp \cdots\] \[\cos(x)=1-\frac{x^2}{2}+\frac{x^4}{4!}\pm\dots\]

OpenStudy (anonymous):

yes.

ganeshie8 (ganeshie8):

then use them

OpenStudy (anonymous):

I've gone the next few steps... but I don't understand how that gives me a zero.!?!?

ganeshie8 (ganeshie8):

\[\large \cos(\sin x) = 1-\dfrac{(\sin x)^2}{2}+\cdots = 1 -\dfrac{(x - \frac{x^3}{3!}+\cdots)^2}{2}+\cdots \] fine with above ?

OpenStudy (anonymous):

yes. got that step. Im afraid I need to be led by the hand still though.

ganeshie8 (ganeshie8):

\[\large{\begin{align} &\color{red}{ \cos(\sin x)}-\color{blue}{\cos x}\\~\\ &= \color{red}{1-\dfrac{(x - \frac{x^3}{3!}+\cdots)^2}{2}+\mathcal{O}(x^4)}-\left(\color{blue}{1-\dfrac{x^2}{2}+\mathcal{O}(x^4)}\right) \\~\\ &= \color{red}{-\dfrac{(x - \frac{x^3}{3!}+\cdots)^2}{2}+\mathcal{O}(x^4)}-\left(\color{blue}{-\dfrac{x^2}{2}+\mathcal{O}(x^4)}\right) \\~\\ &= \color{red}{-\dfrac{x^2+\mathcal{O}(x^4)}{2}+\mathcal{O}(x^4)}-\left(\color{blue}{-\dfrac{x^2}{2}+\mathcal{O}(x^4)}\right) \\~\\ &= \color{red}{-\dfrac{\mathcal{O}(x^4)}{2}+\mathcal{O}(x^4)}-\left(\color{blue}{\mathcal{O}(x^4)}\right) \\~\\ &=\mathcal{O}(x^4) \end{align}}\]

ganeshie8 (ganeshie8):

that shows that the degree of each term in the numerator is at least \(4\)

OpenStudy (anonymous):

I'm so sorry, but whats that o like symbol? starting out with calc

ganeshie8 (ganeshie8):

If it helps, you may replace \(\mathcal{O}(x^4)\) by \(x^4(stuff)\)

OpenStudy (anonymous):

got it. thanks

OpenStudy (anonymous):

this actually pretty clever.

ganeshie8 (ganeshie8):

\[\large{\begin{align} &\color{red}{ \cos(\sin x)}-\color{blue}{\cos x}\\~\\ &= \color{red}{1-\dfrac{(x - \frac{x^3}{3!}+\cdots)^2}{2}+x^4(stuff)}-\left(\color{blue}{1-\dfrac{x^2}{2}+x^4(stuff)}\right) \\~\\ &= \color{red}{-\dfrac{(x - \frac{x^3}{3!}+\cdots)^2}{2}+x^4(stuff)}-\left(\color{blue}{-\dfrac{x^2}{2}+x^4(stuff)}\right) \\~\\ &= \color{red}{-\dfrac{x^2+x^4(stuff)}{2}+x^4(stuff)}-\left(\color{blue}{-\dfrac{x^2}{2}+x^4(stuff)}\right) \\~\\ &= \color{red}{-\dfrac{x^4(stuff)}{2}+x^4(stuff)}-\left(\color{blue}{x^4(stuff)}\right) \\~\\ &=x^4(stuff) \end{align}}\]

OpenStudy (anonymous):

oh no need for that. I worked it out too. thanks a lot lot lot

OpenStudy (astrophysics):



OpenStudy (anonymous):

I'm like that after every single calc question. so many...So many. I'm probably going to keep asking today

ganeshie8 (ganeshie8):